OpenStudy (anonymous):

How to find the number of common odd positive divisors between two numbers? the two numbers are 27900 and 20700. I've put them into their prime factorisation but I can't remember what to do next?

ganeshie8 (ganeshie8):

\(27900 = 2^2\times 3^2 \times 5^2 \times 31\) \(20700 = 2^2 \times 3^2 \times 5^2 \times 23\)

ganeshie8 (ganeshie8):

Since you're interested in odd common divisors, discard the even prime "2"

ganeshie8 (ganeshie8):

do u know number of divisors function ?

OpenStudy (anonymous):

as in Euler's Totient function?

ganeshie8 (ganeshie8):

I think it is represented using \(\large \sigma \)

ganeshie8 (ganeshie8):

let \(N = p_1^{a_1} p_2^{a_2}p_3^{a_3}.... \) be the prime factorization of some number \(N\) then, number of divisors of \(N\) is given by : \(\large \sigma = (a_1+1)(a_2+2)(a_3+3)... \)

ganeshie8 (ganeshie8):

take the GCF of both numbers and find the number of divisors, you're done

OpenStudy (anonymous):

so assumingly if the number of divisors is written in such a way, then there has to be some sort of ordering for example: \[27900=2^2 \times 3^2 \times 5^2 \times 31\] but as you said we discard the 2, how do we know which power, so to speak, to add the 1 to, which to add 2 to, etc...as surely you will get a different value depending which power you take to be \[\alpha_1\] etc I may have just completely overcomplicated that, I haven't done this in a long time

ganeshie8 (ganeshie8):

lets finish solving the problem first, and get back to ur question maybe :)

ganeshie8 (ganeshie8):

\(27900 = 2^2\times 3^2 \times 5^2 \times 31\) \(20700 = 2^2 \times 3^2 \times 5^2 \times 23\) GCF excluding even primes = \( 3^2 \times 5^2\) So, number of odd common divisors = \((2+1)(2+1) = 9\)

OpenStudy (anonymous):

oh I get what you mean now!! Sorry lol

ganeshie8 (ganeshie8):

good, easy to see it for oneself... than explaining :)

OpenStudy (fibonaccichick666):

why is it plus one for both ganeshie?(personal curiosity)

ganeshie8 (ganeshie8):

just combinatorics @FibonacciChick666

OpenStudy (anonymous):

Thank you so much!!

ganeshie8 (ganeshie8):

\(N = p_1^{a_1} p_2^{a_2}p_3^{a_3}.... \) How many ways you can choose \(p_1\) ?

OpenStudy (fibonaccichick666):

ahh, ok I thought based on the phi function that the second would be plus 2. Thanks Ganeshie

OpenStudy (fibonaccichick666):

one way, alpha sub 1 times

ganeshie8 (ganeshie8):

powers of \(p_1\) can be : \(0, 1, 2, ..... a_1 \) that gives \(a_1 + 1\) ways to choose a power of \(p_1\)

OpenStudy (fibonaccichick666):

Ok, but why is \[\large \sigma = (a_1+1)(a_2+2)(a_3+3)...\] Saying to use the progression 1,2,3... ?

ganeshie8 (ganeshie8):

sorry, i made a mistake earlier it should be : \(\large \sigma = (a_1+1)(a_2+1)(a_3+1)... \)

OpenStudy (anonymous):

I see what you mean now, I've just answered my own question trying to give an example haha

ganeshie8 (ganeshie8):

:) thanks for catching the mistake @FibonacciChick666

OpenStudy (ilovecake):

ok so I do not need to help you

OpenStudy (anonymous):

sorry just quick question, given that it specifies for POSITIVE common odd divisors, when would you not get positive divisors? and if you didn't would you then just have to half it?

OpenStudy (fibonaccichick666):

ahh ok, I was confused thanks. That clears it up. @ganeshie8 (I was following this because I just couldn't remember how to solve it. Can't believe I lost that musch number theory in a semester!)

OpenStudy (anonymous):

@FibonacciChick666 I'm exactly the same...I read the question and I was like whaaaaT? haha

OpenStudy (ilovecake):

Bye:) and good luck

ganeshie8 (ganeshie8):

same wid me lol @sarahusher negative divisor = -positive divisor

OpenStudy (fibonaccichick666):

@sarahusher haha yea, it's never fun to forget a class you need which was difficult to learn in the first place, but Ganeshie has ya covered.

ganeshie8 (ganeshie8):

so if u want all the divisors, just take twice of positive divisors

OpenStudy (anonymous):

and for just the total number of common positive divisors, would I do the same but leave the 2 in there too!

OpenStudy (fibonaccichick666):

yup

OpenStudy (ilovecake):

Do you want me to help or NOT

ganeshie8 (ganeshie8):

yup take it on the complete factorization

OpenStudy (anonymous):

so in this case the total number of common divisors will be 27?

OpenStudy (ilovecake):

ahhhhhhhhhhh just forget it!!

ganeshie8 (ganeshie8):

\(27900 = 2^2\times 3^2 \times 5^2 \times 31\) \(20700 = 2^2 \times 3^2 \times 5^2 \times 23\) GCF = \( 2^2\times3^2 \times 5^2\) # of common divisors = \((2+1)(2+1)(2+1)\)

OpenStudy (anonymous):

great, thank you for clearing this up for me!! Helped a lot!! :-)

ganeshie8 (ganeshie8):

np :)